OpenStudy (anonymous):

If the instructions for a problem ask you to use the smallest possible domain to completely graph two periods of y = 5 + 3 cos 2(x - pi/3), what should be used for Xmin and Xmax? Explain your answer.

OpenStudy (anonymous):

i would say the following two periods could be from \(0\) to \(4\pi\)

OpenStudy (anonymous):

set \(2x-\frac{\pi}{3}=0\) and set \[2x-\frac{\pi}{3}=2\pi\] and solve for \(x\)

OpenStudy (anonymous):

do I not need the "y=5+3" part? @satellite73

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh i made a typo though damn

OpenStudy (anonymous):

set \[2x-\frac{\pi}{3}=0\] and set \[2x-\frac{\pi}{3}=4\pi\] since you want two full periods

OpenStudy (anonymous):

of course you could also use \[2x-\frac{\pi}{3}=-2\pi\] and \[2x-\frac{\pi}{3}=2\pi\] makes no difference

OpenStudy (anonymous):

okay! then just solve for x correct? thanks! I just wasnt sure how to "explain my answer" lol I mean I graphed it but wasnt sure! Thanks!

OpenStudy (anonymous):

yw